3) A consumers utility function is u(r,y (a) Find the consumers optimal choice for x, y as functions of income I and Be careful! (b) Sketch the demand curves for r,y as functions of income I when prices are p 16,py 2. (Be careful!)

Textbook: Nicholson & Snyder, Microeconomic Theory, 12th edition.
Answer #1

A) MRS = MUx/ MUy

MUx = 1

MUy = dU/dY = (-1/2)*-2*y​​​​​​-3

MUy = 1/y​​​​​​3

MRS= y​​​​3

so at eqm , MRS = Px/Py

y​​​​​​3​​​​​= Px/Py

so y* = (Px/Py)1/3

Now from BC:

XPx + YPy = I

So X = (I-YPy)/Px

X* = [I - (Px/Py)1/3 *Py]/Px

X* = [I - Px​​​1/3Py2/3]/Px

B) Y*= (16/2)1/3 = 2

X* = [ I - 2*2]/16

X* = (I-4)/16

graph

Y* is independent of income

For X:

Income X

​​​​​​